# LDP for invariant measures of the Allen-Cahn equation with inhomogeneous boundary conditions and multiplicative noise

### v1

@@ 4.3 Upper bound
Assume the claim does not hold. Then, there exist $\delta>0$, $r>0$ and a sequence of function $z_{n}\in C([-T_{n},0];E)$ with $T_{n}\nearrow+\infty$ such that $\beta_{n}:=\|z(-T_{n})-(m_{L}-\psi)\|_{E}\searrow 0$ and
@@
and $C_{r}$ is a constant depends on $r$. As the level sets of $\mathbf{E}^{*}_{L}(\cdot)$ are compact in $E$, we have that $K_{\Lambda,-k,0}$ are compact in C([k,0];E)C([-k,0];E). By taking $k=1$ we can find $\{z_{n_{1}}\}\subseteq\{z_{n}\}$ and $\hat{z}_{1}\in C([-1,0];E)$ such that $z_{n_{1}|_{[-1,0]}}$ converges to $\hat{z}_{1}$ in C([1,0];E)C([-1,0];E). With the same arguments, we can find a subsequence $\{z_{n_{2}}\}\subset\{z_{n_{1}}\}$ and $\hat{z}_{2}\in C([-2,0];E)$ such that $z_{n_{2}|_{[-2,0]}}$ converges to $\hat{z}_{2}$ in C([2,0];E)C([-2,0];E). Continuing this procedure, we can find a subsequence $\{z_{n^{\prime}}\}\subset\{z_{n}\}$ converging to some $\hat{z}$ in $C((-\infty,0];E)$. Thanks to Theorem 5.1 of [cerrai2004large], $I_{-k}(\cdot)$ is lower semi-continuous so we have $I_{-k}(\hat{z})\leq r$ for any $k\geq 0$ which implies $I_{-\infty}(\hat{z})\leq r$. Moreover, as $\beta_{n}\searrow 0$, we have
@@
Therefore, $\hat{z}\in K_{-\infty}(r)$. Thus, thanks to the characterization of $U_{L}$ (see Proposition 3.8), $U_{L}(\hat{z}(0))\leq r$ and $\hat{z}(0)\in\Phi_{L}(r)$. However, by our construction
@@
Assume there exist $r>0$, $\delta>0$ and $\rho>0$ such that for any $n\in\mathbb{N}$ there exists $u\in H_{\rho,r,\delta}(n)$ with
@@
Then, by Proposition 3.2 and (3.12), $\mathbf{E}^{*}_{L}(u(n))\leq C_{r,\rho}$, for all $n\in\mathbb{N}$. Here, $C_{r,\rho}$ is positive constant depending on $r$ and $\rho$. Now, for any $k\in\mathbb{N}$ we define
@@
By same arguments as in the proof of Lemma 3.7, for any $\eta>0$, there exist $\bar{k}$ sufficiently large, $N\in\mathbb{N}$ and some $t\in[\frac{\bar{k}}{2},\bar{k}]$ such that for any $i\in\mathbb{N}$, $i\geq N$
@@
Thus, by FWULDP results, there exists $\epsilon_{0,j}>0$ such that for any $\epsilon\leq\epsilon_{0,j}$,
@@
Under Assumption 1, there exists constant $C>0$ and $\delta_{0}>0$ such that for every $0<\delta<\delta_{0}$ there exists $\epsilon_{0}>0$ such that
@@
In particular, the measures $\mu_{\epsilon}$ concentrate around the minimizer $m_{L}$ exponentially fast, when $\epsilon$ is sufficiently small.

### v2

@@ 4.3 The LDP upper bound
Assume the claim does not hold. Then, there exist $\delta>0$, $r>0$ and a sequence of functions $z_{n}\in C([-T_{n},0];E)$ with $T_{n}\nearrow+\infty$ such that $\beta_{n}:=\|z(-T_{n})-(m_{L}-\psi)\|_{E}\searrow 0$ and
@@
and $C_{r}$ is a constant depending on $r$. As the level sets of $\mathbf{E}^{*}_{L}(\cdot)$ are compact in $E$, we find that $K_{\Lambda,-k,0}$ are compact in C([k,0];E)C([-k,0];E). Taking $k=1$, we can find $\{z_{n_{1}}\}\subseteq\{z_{n}\}$ and $\hat{z}_{1}\in C([-1,0];E)$ such that $z_{n_{1}|_{[-1,0]}}$ converges to $\hat{z}_{1}$ in C([1,0];E)C([-1,0];E). With the same arguments, we can find a subsequence $\{z_{n_{2}}\}\subset\{z_{n_{1}}\}$ and $\hat{z}_{2}\in C([-2,0];E)$ such that $z_{n_{2}|_{[-2,0]}}$ converges to $\hat{z}_{2}$ in C([2,0];E)C([-2,0];E). Continuing this procedure, we can find a subsequence $\{z_{n^{\prime}}\}\subset\{z_{n}\}$ converging to some $\hat{z}$ in $C((-\infty,0];E)$. Thanks to Theorem 5.1 of [cerrai2004large], $I_{-k}(\cdot)$ is lower semi-continuous, so we have $I_{-k}(\hat{z})\leq r$ for any $k\geq 0$ which implies $I_{-\infty}(\hat{z})\leq r$. Moreover, as $\beta_{n}\searrow 0$, we have
@@
Therefore, $\hat{z}\in K_{-\infty}(r)$. Thus, by the characterisation of $U_{L}$ (see Proposition 3.8), $U_{L}(\hat{z}(0))\leq r$ and $\hat{z}(0)\in\Phi_{L}(r)$. However, by our construction
@@
Assume there exist $r>0$, $\delta>0$, and $\rho>0$ such that for any $n\in\mathbb{N}$ there exists $u\in H_{\rho,r,\delta}(n)$ with
@@
Then, by Proposition 3.2 and (3.12), $\mathbf{E}^{*}_{L}(u(n))\leq C_{r,\rho}$, for all $n\in\mathbb{N}$. Here, $C_{r,\rho}$ is a positive constant depending on $r$ and $\rho$. Now, for any $k\in\mathbb{N}$ we define
@@
By the same arguments as in the proof of Lemma 3.7, for any $\eta>0$, there exist $\bar{k}$ sufficiently large, $N\in\mathbb{N}$ and some $t\in[\frac{\bar{k}}{2},\bar{k}]$ such that for any $i\in\mathbb{N}$, $i\geq N$
@@
Thus, according the FWULDP results, there exists $\epsilon_{0,j}>0$ such that for any $\epsilon\leq\epsilon_{0,j}$,
@@
Under Assumption 1, there exist $C>0$ and $\delta_{0}>0$ such that for every $0<\delta<\delta_{0}$ there exists $\epsilon_{0}>0$ such that
@@
In particular, the measures $\mu_{\epsilon}$ concentrate around the minimiser $m_{L}$ exponentially fast, when $\epsilon$ is sufficiently small.